# Preparations for Quantum Computing in Hadron Physics

### v1

@@ Abstract
Quantum computers are coming online and will quickly impact hadron physics once certain fidelity, decoherence and memory thresholds are met, quite possibly within a decade. We review a selected number of topics where ab-initio QCD-level information about hadrons can be obtained with this computational tool that is hard to come by from other methods. This includes high baryon-density systems such as neutron-star matter (with a sign problem in lattice gauge theory); fragmentation functions; Monte Carlo generation of particles which accounts for quantum correlations in the final state; entropy production in jets; and generally, any application where time evolution in Minkowski space (as opposed to a Euclidean formulation) or where large chemical potentials play an important dynamical role. For other problems, such as the prediction of very highly excited hadron spectroscopy, they will not be a unique, but a complementary tool.
Keywords: Hadron Physics; Quantum computing; Equation of State; Neutron Stars; Fragmentation Functions; Monte Carlo simulations; Evolution.
@@ 1 Introduction: Strokebrushes of Hadron Physics circa 2025
The hadron spectrum has grown to hundreds of particles and resonances by now. The basic mesons and baryons are all reasonably predicted by Monte Carlo methods in the Lattice Gauge Theory (LGT) formulation, after forty years of improvements [2, 4, 3]. The current thrust of experiment and phenomenology is in sorting out the excited spectrum above strong-force decay thresholds. This includes possible tetra/pentaquark states, also among them multiheavy hadrons (containing several heavy quarks) [5, 6]. LGT practitionners [7] are also developing techniques to confront this excited spectrum, learning how to embed states in a continuum of several ground-state hadrons, and because such difficulties mostly stem from the discretization of spacetime, any computer calculation, including those in quantum computers, will also have to face them.
The same applies to the Lattice’s difficulties with chiral fermions: the Nielsen-Ninomiya theorem [8] applies equally to LGT whether formulated in Euclidean space for a classical computer or on Minkowski space for a quantum one. Thus, the quantum hardware with spacetime discretization is not expected to be advantageous to deal with problems related to chiral symmetry and its breaking, such as the insensitivity (or sometimes “restoration”, a less appropriate term) of the higher spectrum to chiral symmetry breaking [9] bringing about parity doublets. In section 5 below we discuss a particle–based encoding for quantum computers which might avoid some of these problems.
The one important question which really remains unanswered in the hadron spectrum is Nathan Isgur’s Where is the glue[10]? Although hybrid mesons [11] and glueballs [12] have long been computed in LGT [13, 14, 15], their experimental extraction is, to say it candidly, stuck. Glueballs are mostly mixed with quarkonia into the meson spectrum; $J^{PC}$ exotics which could be hybrid-meson candidates do not abound, and the only broadly accepted one [16], the $1^{-+}$ $\pi_{1}(1600)$ is somewhat too light to perfectly fit the tab as the expected exotic hybrid. In any case, it is unclear what contribution quantum computers could make here either: conventional theoretical techniques give reasonable estimates of the parameters and decay channels of hybrid mesons, with incremental but sure improvement. This does not mean that quantum computers should not be deployed to calculate the hadron spectrum [17, 18], and indeed heavy mesons and all-heavy baryons in the Cornell model have already been recalculated employing the quantum computer as a small diagonalizer, but this should be seen as a means to calibrate the quality of their computations, not for the product of these.
@@
Having discussed the modest interest of employing quantum computers in hadron spectroscopy, let us examine their possible role in advancing hadron structure and reactions. Elastic and few-channel inelastic hadron-hadron interactions are well understood at low energies by a combination of methods: chiral effective theory [20] and dispersion relations [21]. These are clever parametrizations of the experimental data that allow some level of control of the systematic uncertainties [22], but have more parameter freedom than the very stringent number of parameters in the Lagrangian of QCD allows.
@@
Next, let us briefly discuss the collective medium formed in heavy ion collisions. A new phase of matter has been found in these, which behaves as a fluid [27], not a rarefied gas of particles. Near the phase transition to a more conventional hadron gas, it has low viscosity [28], relative to its entropy density. We hesitate it to call it, as is common, a “Quark–Gluon Plasma” [29] because of this strongly-coupled fluid behaviour as well as the predicted appearance of chains or loops of confined colour charges, which would make its interpretation as a plasma very difficult, so a “Quark–Gluon Gasoline” would be a closer analogy at the temperatures of order 150 MeV near the phase transition [30, 31]. Be it as it may, Lattice Gauge Theory [32] is a tool which can calculate many properties of hadron matter at very large temperatures, as long as the baryon chemical potential is negligible (see Section 2 shortly).
Many more interesting topics have developed within hadron physics in the last years. For example, the so called “Strong CP problem” seems to have been put to rest [33, 34] by showing that the $\theta$ parameter does not show in CP-violating observable, if the limits of large volume and sum over topological sectors are taken in this order, unlike in earlier work which produced a CP-violating phase from the $\theta$-vacuum. This is the realm of analytical work.
Also, artificial intelligence is making inroads in our field[35, 36], to assist event or jet classification, or identifying parameters of scattering amplitudes such as pole positions, for example. All these examples, and many more, show the richness of physics and methods in hadron physics: many of them will not see substantial progress due to the arrival of quantum computers. However, there are a few interesting open problems where they will have an impact, and we devote the next section 2 to list three of them. Additional material, with overlaps to nuclear physics and to particle physics, can be found in the white paper [37] for the US DOE.
2 Some niche problems which remain unaddressed from first principles
@@ 2.1 Equation of state of neutron stars
The behaviour of an atomic nucleus composed by a few nucleons [38] is well characterized by chiral interactions. Nuclear matter below the saturation density $n_{s}$ may be addressed from the same chiral interactions or even directly from scattering data without assuming a Hamiltonian [39]. But as density increases, probably around 1.5-2 times $n_{s}$, certainly above 2-2.5 $n_{s}$, the chiral expansion breaks down and uncertainties become unmanageable [40].
From there on, the allowed band of Equations of State inside a neutron star is constrained only by fundamental theory, namely by causality, $c_{s}^{2}\leq c^{2}=1$ and mechanical stability, $c_{s}^{2}\geq 0$; since $c_{s}^{2}=\frac{dP}{d\varepsilon}$ con $P$ the pressure and $\varepsilon$ the mass-energy density. This is seen in figure 2.
@@
Narrowing the band of possible equations of state in figure (2) as much as possible is important to potentially run tests of the theory of gravity with good knowledge, from microscopic physics alone, of the stress-energy tensor, the right hand side of Einstein’s field equations (for an isotropic and homogeneous fluid at rest, $T={\rm diag}(\varepsilon,P,P,P)$):
@@
One possibility is to work from the nuclear side (from left to center of the figure) by employing the quantum computer, basically, as a diagonalizer, profiting from its large (variational) Hilbert space, with basis of size $N=2^{n_{q}}$ exponentiating the number of qubits $n_{q}$. This method would not suffer from the breakdown of the perturbative expansion in Chiral Perturbation Theory upon computing thermodynamic properties. However it would still be subject to the failure of the theory in providing an ab initio Hamiltonian as its organizing principle, that Chiral Expansion in powers of momentum etc. would fail; even the degrees of freedom, nucleons, would be in doubt at such large densities than the vacuum symmetries [45, 46] (from which the Lagrangian is constructed) could change.
@@
2.2 Fragmentation functions in hadronizing jets
@@
Not such progress has been made, however, in computing fragmentation functions $D(x)$. These are nonetheless extremely interesting objects because (a) they grant access to all hadrons which can be reconstructed in the final state, and not only the much smaller number of them which are stable enough to be prepared for the experiment in the initial state; and (b) they must be sensitive to how quarks and gluons confine into hadrons, the asymptotic end states in any hadron process (a suggestive image for this jet hadronization is to think of a nut, bolt or spring moving at large speed, picking up other pieces as it propagates, and producing an entire mechanical clock in the end state). We do not have so many experimental windows to actual confinement mechanisms such as the center-vortex one, which remains mostly a theoretical endeavour.
A possible definition of fragmentation functions employs the probability of finding a hadron $h$ plus other particles $X_{\text{out}}$ not reconstructed, within the jet that was launched by a given bare quark or gluon $j$ whose longitudinal momentum fraction $z=p_{h}^{-}/p_{q}^{-}$ is the argument of the fragmentation function [48, 49, 50] $D^{h}_{j}(z)$, which at given renormalization scale and on a discretized momentum space is defined by
$D_{j}^{h}(z)\equiv\frac{\rm Tr_{c}}{N_{c,j}}\sum_{X}\langle j,p_{1}|h,X_{\text{out}}\rangle\langle h,X_{\text{out}}|j,p_{1}\rangle$ (2)
with $N_{c,j}$ being the number of colors, and the trace also being applied over parton colour space.
@@
Euclidean LGT sztumbles because the fields are taken at different points along light–front time (note the integral over $x^{+}$), an evolution operator along a lightlike direction is needed (which is not possible in an Euclidean formulation with four equivalent spacelike directions).
Now, near-to-medium future quantum computers promise quick improvement. Currently, Nambu-Jona-Lasinio modelcomputations are already at hand [75], and our group is actively exploring the hardware needed to proceed to appropriately regulated Quantum Chromodynamics, with its more fundamental Hamiltonian formulated in Light-Front quantization (see subsec. 6.2 and section 8 below).
@@ 2.3 Monte Carlo event generation
Another interesting example which affects many experimental collaborations in high-energy physics is particle production in Monte Carlo generators. The best known example is PYTHIA in which from the initial hard partons, strings with quark/antiquark sources are generated that then produce the final state hadrons. This production process follows a classical probability distribution. It has been very successful, for decades now, in giving a rough parametrization of one–body spectra (with various parameter “tunes”) and is widely appreciated by its versatility and speed in producing numerous multiparticle events.
However, in the last years, the flourishing experimental works on particle-particle two-body correlations have stressed the program. For example, the ALICE collaboration [53] reported a qualitative disagreement between their data, shown in figure 4 and that shows baryon-baryon anticorrelation (if a baryon is emitted in a given azimuthal direction, it is less likely than a second baryon is emitted along a nearby direction), unlike mesons, whereas the Monte Carlo simulation yields a positive correlation analogous to that of mesons (which are bosons and are likely to be closely emitted in momentum space).
Figure 4: From left to right: a) Typical ALICE [53] experimental data and Monte Carlo simulation of two-meson correlation against $\varphi$, with a peak at $\phi_{1}-\phi_{2}=0$ indicating that the mesons are positively correlated. b) and c) In the experiment, the baryon-baryon and antibaryon-antibaryon correlations dip below 1, implying anticorrelation. In this baryon case, Monte Carlo simulations disagree with the experimental data. This is true for both identical and different baryon species. Copied from [53] by the ALICE collaboration under the terms of the Creative Commons License 4.0 (http://creativecommons.org/licenses/by/4.0/); no changes have been effected.
A possible ad-hoc fix [54] (see [55] for further discussion) exemplifying the root of the problem in the Monte Carlo simulation is to spread the baryons in phase space, with two rules “One-baryon policy” and “All baryon policy” which force every string to produce one and only one baryon (the first rule to guarantee the anticorrelation, the second to avoid lessening the overall number of baryons).
@@
A much more satisfactory solution would be to have a full quantum calculation where the Pauli principle among quarks is satisfied and where correlations and anticorrelations satisfy the same quantum rules that apply in the real world. This is however computationally unfeasible due to the huge running times: a quantum computer, however, could probably speed-up the hadronization steps and produce more natural (anti)correlations among particles.
@@ 3.1 Digital versus Analog Simulations
Digital simulations[56] of time evolution in a quantum system (such as a hadron) are based on the application of stroboscopic unitary evolution, first proposed by Lloyd [57], on universal quantum computers. If a Hamiltonian consists of the sum of several terms $H=\sum_{j}H_{j}$, a Suzuki-Trotter type formula with stroboscopic (slow-motion) steps is employed,
$e^{-iHt}\simeq\left(\prod_{j}e^{-iH_{j}\frac{t}{n}}\right)^{n}$ (5)
@@
In contrast, analog quantum computations are based on the fact that there are systems, the simulators, that can be tuned to behave in certain regions of parameter space as another system which is in principle inaccessible[56]. The mapping should relate the states of both systems, if $\left|\phi_{t}\right\rangle$ is an state of the objective system and $\left|\psi_{t}\right\rangle$ the corresponding state on the simulator, there should exist a bijection such that
@@
In general analog simulators are robust against errors, at least more than their digital counterparts and there is no need of Trotter decompositions to implement unitary evolution.
Both digital and analog quantum computers can be used to execute quantum algorithms. The work of Jordan, Lee, and Preskill[58] about real-time scattering in $\phi^{4}$ theory was one of the earliest algorithms in the context of high-energy physics. The simulation of a scattering event was structured into three stages: a) preparation of the asymptotic wavefunction, b) time evolution through the scattering region, and c) measurement of the outgoing states. These stages mirror the fundamental routines of every quantum computation: state preparation, unitary evolution, and measurement and can be executed in both types of computers. The asymptotic costs of each step were also calculated, demonstrating that the process can be efficiently simulated. Numerous other theoretical studies followed on Lattice Gauge Theories [62, 59, 60, 63, 64, 65, 61], spectroscopy[17, 18], jet evolution[66, 67, 68], fragmentation[75], encoding [70, 69, 71], and more [72]; and experimental tests, see for example [73, 74, 76, 77, 78, 79, 80] for digital computations and [81, 82, 83, 84, 85] for analog experiments. The reader is referred to the subsections below for recent applications to hadron physics.

### v2

@@ Abstract
Quantum computers are coming online and will quickly impact hadron physics once certain fidelity, decoherence and memory thresholds are met, quite possibly within a decade. We review a selected number of topics where ab-initio Quantum Chromodynamics-level information about hadrons can be obtained with this computational tool that is hard to come by from other methods. This includes high baryon-density systems such as neutron-star matter (with a sign problem in lattice gauge theory); fragmentation functions; Monte Carlo generation of particles which accounts for quantum correlations in the final state; entropy production in jets; and generally, any application where time evolution in Minkowski space (as opposed to a Euclidean formulation) or where large chemical potentials play an important dynamical role. For other problems, such as the prediction of very highly excited hadron spectroscopy, they will not be a unique, but a complementary tool.
keywords:
Hadron Physics; Quantum computing; Equation of State; Neutron Stars; Fragmentation Functions; Monte Carlo simulations; Evolution.
@@ 1 Introduction: Strokebrushes of Hadron Physics circa 2025
The hadron spectrum has grown to hundreds of particles and resonances by now. The basic mesons and baryons are all reasonably predicted by Monte Carlo methods in the Lattice Gauge Theory (LGT) formulation, after forty years of improvements [2, 4, 3]. The current thrust of experiment and phenomenology is in sorting out the excited spectrum above strong-force decay thresholds. This includes possible tetra/pentaquark states, also among them multiheavy hadrons (containing several heavy quarks) [5, 6]. LGT practitioners [7] are also developing techniques to confront this excited spectrum, learning how to embed states in a continuum of several ground-state hadrons, and because such difficulties mostly stem from the discretisation of spacetime, any computer calculation, including those in quantum computers, will also have to face them.
The same applies to the Lattice’s difficulties with chiral fermions: the Nielsen-Ninomiya theorem [8] applies equally to LGT whether formulated in Euclidean space for a classical computer or on Minkowski space for a quantum one. Thus, the quantum hardware with spacetime discretisation is not expected to be advantageous to deal with problems related to chiral symmetry and its breaking, such as the insensitivity (or sometimes “restoration”, a less appropriate term) of the higher spectrum to chiral symmetry breaking [9] bringing about parity doublets. In section 5 below we discuss a particle–based encoding for quantum computers which might avoid some of these problems.
The one important question which really remains unanswered in the hadron spectrum is Nathan Isgur’s Where is the glue[10]? Although hybrid mesons [11] and glueballs [12] have long been computed in LGT [13, 14, 15], their experimental extraction is, to say it candidly, stuck. Glueballs are mostly mixed with quarkonia into the meson spectrum; $J^{PC}$ exotics which could be hybrid-meson candidates do not abound, and the only broadly accepted one [16], the $1^{-+}$ $\pi_{1}(1600)$ is somewhat too light to perfectly fit the bill as the expected exotic hybrid. In any case, it is unclear what contribution quantum computers could make here either: conventional theoretical techniques give reasonable estimates of the parameters and decay channels of hybrid mesons, with incremental but sure improvement. This does not mean that quantum computers should not be deployed to calculate the hadron spectrum [17, 18], and indeed heavy mesons and all-heavy baryons in the Cornell model have already been recalculated employing the quantum computer as a small diagonaliser, but this should be seen as a means to calibrate the quality of their computations, not for the product of these.
@@
Having discussed the modest interest of employing quantum computers in hadron spectroscopy, let us examine their possible role in advancing hadron structure and reactions. Elastic and few-channel inelastic hadron-hadron interactions are well understood at low energies by a combination of methods: chiral effective theory [20] and dispersion relations [21]. These are clever parametrisations of the experimental data that allow some level of control of the systematic uncertainties [22], but have more parameter freedom than the very stringent number of parameters in the Lagrangian of QCD allows.
@@
Next, let us briefly discuss the collective medium formed in heavy ion collisions. A new phase of matter has been found in these, which behaves as a fluid [27], not a rarefied gas of particles. Near the phase transition to a more conventional hadron gas, it has low viscosity [28], relative to its entropy density. We hesitate to call it, as is common, a “Quark–Gluon Plasma” [29] because of this strongly-coupled fluid behaviour as well as the predicted appearance of chains or loops of confined colour charges, which would make its interpretation as a plasma very difficult, so a “Quark–Gluon Gasoline” would be a closer analogy at the temperatures of order 150 MeV near the phase transition [30, 31]. Be it as it may, Lattice Gauge Theory [32] is a tool which can calculate many properties of hadron matter at very large temperatures, as long as the baryon chemical potential is negligible (see Section 2 shortly).
Many more interesting topics have developed within hadron physics in the last years. For example, the so called “Strong CP problem” seems to have been put to rest [33, 34] by showing that the $\theta$ parameter does not show in CP-violating observables, if the limits of large volume and sum over topological sectors are taken in this order, unlike in earlier work which produced a CP-violating phase from the $\theta$-vacuum. This is the realm of analytical work.
Also, artificial intelligence is making inroads in our field[35, 36], to assist event or jet classification, or identifying parameters of scattering amplitudes such as pole positions, for example. All these examples, and many more, show the richness of physics and methods in hadron physics: many of them will not see substantial progress due to the arrival of quantum computers. However, there are a few interesting open problems where they will have an impact, and we devote the next section 2 to list three of them. It is our intent in this work to confine ourselves to topics in which individual hadrons, their excitations, and their quark-gluon components come to the fore, avoiding excursions into either High Energy Physics or Nuclear Physics, where other reviews have already appeared. Additional material, with overlaps to nuclear physics and to particle physics, can be found, for example, in the white paper [37] for the US DOE.
2 Some niche problems which remain unaddressed with first principles
@@ 2.1 Equation of state of neutron stars
The behaviour of an atomic nucleus composed of a few nucleons [38] is well characterised by chiral interactions. Nuclear matter below the saturation density $n_{s}$ may be addressed with the same chiral interactions or even directly from scattering data without assuming a Hamiltonian [39]. But as density increases, probably around 1.5-2 times $n_{s}$, certainly above 2-2.5 $n_{s}$, the chiral expansion breaks down and uncertainties become unmanageable [40].
From there on, the allowed band of Equations of State inside a neutron star is constrained only by fundamental theory, namely by causality, $c_{s}^{2}\leq c^{2}=1$ and mechanical stability, $c_{s}^{2}\geq 0$; since $c_{s}^{2}=\frac{dP}{d\varepsilon}$ with $P$ the pressure and $\varepsilon$ the mass-energy density. This is seen in figure 2.
@@
Narrowing the band of possible equations of state in figure 2 as much as possible is important to potentially run tests of the theory of gravity with good knowledge, from microscopic physics alone, of the stress-energy tensor, the right hand side of Einstein’s field equations (for an isotropic and homogeneous fluid at rest, $T={\rm diag}(\varepsilon,P,P,P)$):
@@
One possibility is to work from the nuclear side (from left to centre of the figure) by employing the quantum computer, basically, as a diagonaliser, profiting from its large (variational) Hilbert space, with basis of size $N=2^{n_{q}}$ exponentiating the number of qubits $n_{q}$. This method would not suffer from the breakdown of the perturbative expansion in Chiral Perturbation Theory upon computing thermodynamic properties. However it would still be subject to the failure of the theory in providing an ab initio Hamiltonian as its organising principle, that Chiral Expansion in powers of momentum etc. would fail; even the degrees of freedom, nucleons, would be in doubt at such large densities that the vacuum symmetries [45, 46] (from which the Lagrangian is constructed) could change.
@@
2.2 Fragmentation functions in hadronising jets
@@
Not such progress has been made, however, in computing fragmentation functions $D(x)$. These are nonetheless extremely interesting objects because (a) they grant access to all hadrons which can be reconstructed in the final state, and not only the much smaller number of them which are stable enough to be prepared for the experiment in the initial state; and (b) they must be sensitive to how quarks and gluons confine into hadrons, the asymptotic end states in any hadron process (a suggestive image for this jet hadronisation is to think of a nut, bolt or spring moving at large speed, picking up other pieces as it propagates, and producing an entire mechanical clock in the end state). We do not have so many experimental windows to actual confinement mechanisms such as the centre-vortex one, which remains mostly a theoretical endeavour.
A possible definition of fragmentation functions employs the probability of finding a hadron $h$ plus other particles $X_{\text{out}}$ not reconstructed, within the jet that was launched by a given bare quark or gluon $j$ whose longitudinal momentum fraction $z=p_{h}^{-}/p_{q}^{-}$ is the argument of the fragmentation function [48, 49, 50] $D^{h}_{j}(z)$, which at given renormalisation scale and on a discretised momentum space is defined by
$D_{j}^{h}(z)\equiv\frac{\rm Tr_{c}}{N_{c,j}}\sum_{X}\langle j,p_{1}|h,X_{\text{out}}\rangle\langle h,X_{\text{out}}|j,p_{1}\rangle,$ (2)
with $N_{c,j}$ being the number of colours, and the trace also being applied over parton colour space.
@@
Euclidean LGT stumbles because the fields are taken at different points along light–front time (note the integral over $x^{+}$), an evolution operator along a lightlike direction is needed (which is not possible in an Euclidean formulation with four equivalent spacelike directions).
Now, near-to-medium future quantum computers promise quick improvement. Currently, Nambu-Jona-Lasinio model computations are already at hand [75], and our group is actively exploring the hardware needed to proceed to appropriately regulated Quantum Chromodynamics, with its more fundamental Hamiltonian formulated in Light-Front quantisation (see subsec. 6.2 and section 8 below).
@@ 2.3 Monte Carlo event generation
Another interesting example which affects many experimental collaborations in high-energy physics is particle production in Monte Carlo generators. The best known example is PYTHIA in which from the initial hard partons, strings with quark/antiquark sources are generated that then produce the final state hadrons. This production process follows a classical probability distribution. It has been very successful, for decades now, in giving a rough parametrisation of one–body spectra (with various parameter “tunes”) and is widely appreciated by its versatility and speed in producing numerous multiparticle events.
However, in the last years, the flourishing experimental works on particle-particle two-body correlations have stressed the program. For example, the ALICE collaboration [53] reported a qualitative disagreement between their data, shown in figure 4 and that shows baryon-baryon anti-correlation (if a baryon is emitted in a given azimuthal direction, it is less likely than a second baryon is emitted along a nearby direction), unlike mesons whereas the Monte Carlo simulation yields a positive correlation analogous to that of mesons (which are bosons and are likely to be closely emitted in momentum space).
Figure 4: From left to right: a) Typical ALICE [53] experimental data and Monte Carlo simulation of two-meson correlation against $\varphi$, with a peak at $\phi_{1}-\phi_{2}=0$ indicating that the mesons are positively correlated. b) and c) In the experiment, the baryon-baryon and antibaryon-antibaryon correlations dip below 1, implying anti-correlation. In this baryon case, Monte Carlo simulations disagree with the experimental data. This is true for both identical and different baryon species. Copied from [53] by the ALICE collaboration under the terms of the Creative Commons License 4.0 (http://creativecommons.org/licenses/by/4.0/); no changes have been effected.
A possible ad-hoc fix [54] (see [55] for further discussion) exemplifying the root of the problem in the Monte Carlo simulation, is to spread the baryons in phase space, with two rules “One-baryon policy” and “All baryon policy” which force every string to produce one and only one baryon (the first rule to guarantee the anti-correlation, the second to avoid lessening the overall number of baryons).
@@
A much more satisfactory solution would be to have a full quantum calculation where the Pauli principle among quarks is satisfied and where correlations and anti-correlations satisfy the same quantum rules that apply in the real world. This is however computationally unfeasible due to the huge running times: a quantum computer, however, could probably speed-up the hadronisation steps and produce more natural (anti)correlations among particles.
@@ 3.1 Digital versus Analog Simulations
Digital simulations[56] of time evolution in a quantum system (such as a hadron) are based on the application of stroboscopic unitary evolution, first proposed by Lloyd [57], on universal quantum computers. If a Hamiltonian consists of the sum of several terms $H=\sum_{j}H_{j}$, a Suzuki-Trotter type formula with stroboscopic (slow-motion) steps is employed
$e^{-iHt}\simeq\left(\prod_{j}e^{-iH_{j}\frac{t}{n}}\right)^{n},$ (5)
@@
In contrast, analogue quantum computations are based on the fact that there are systems, the simulators, that can be tuned to behave in certain regions of parameter space as another system which is in principle inaccessible[56]. The mapping should relate the states of both systems, if $\left|\phi_{t}\right\rangle$ is an state of the objective system and $\left|\psi_{t}\right\rangle$ the corresponding state on the simulator, there should exist a bijection such that
@@
In general analogue simulators are robust against errors, at least more than their digital counterparts and there is no need of Trotter decompositions to implement unitary evolution.
Both digital and analogue quantum computers can be used to execute quantum algorithms. The work of Jordan, Lee, and Preskill[58] about real-time scattering in $\phi^{4}$ theory was one of the earliest algorithms in the context of high-energy physics. The simulation of a scattering event was structured into three stages: a) preparation of the asymptotic wavefunction, b) time evolution through the scattering region, and c) measurement of the outgoing states. These stages mirror the fundamental routines of every quantum computation: state preparation, unitary evolution, and measurement and can be executed in both types of computers. The asymptotic costs of each step were also calculated, demonstrating that the process can be efficiently simulated. Numerous other theoretical studies followed on Lattice Gauge Theories [62, 59, 60, 63, 64, 65, 61], spectroscopy[17, 18], jet evolution[66, 67, 68], fragmentation[75], encoding [70, 69, 71], and more [72]; and experimental tests, see for example [73, 74, 76, 77, 78, 79, 80] for digital computations and [81, 82, 83, 84, 85] for analogue experiments. The reader is referred to the subsections below for recent applications to hadron physics.